# Constrained and Ordered Level Planarity Parameterized by the Number of Levels

### v1

@@ Abstract
keywords:
Parameterized Complexity, Graph Drawing, XNLP, XP, W[t]-hard, Level Planarity, Planar Poset Diagram, Computational Geometry
@@ 1 Introduction
Klemz and Rote [24] showed that OLP (and, thus, CLP) is NP-hard even when restricted to the case where the underlying undirected graph of $G$ is a disjoint union of paths. Note that such a graph has bounded pathwidth (and treewidth), maximum degree, and feedback vertex set number, ruling out efficient parameterized algorithms with respect to these classical parameters. Additionally, the instances produced by their reduction have a levelwidth of only two. Independently, Brückner and Rutter [7] provided a very different reduction for showing that CLP is NP-hard (in fact, their reduction shows the NP-hardness of Partial Level Planarity, which can be seen as a generalization of OLP and a special case of CLP; see below). The instances constructed by their reduction are connected and have bounded maximum degree. They also present a polynomial time algorithm for CLP for the case where the graph has a single sink, which they later sped up [8].
@@
As stated above, Klemz and Rote [24] showed that OLP (and, thus, CLP) is NP-hard even when restricted to instances of levelwidth two. In this paper, we focus on the other “dimension”: we study CLP and OLP parameterized by height.
@@
Parameterizing OLP by height captures the “linear” nature of the solution – this is reminiscent of recent results by Bodlaender, Groenland, Jacob, Jaffke, and Lima [5] who established XNLP-completeness for several problems parameterized by linear width measures (e.g., Capacitated Dominating Set by pathwidth and Max Cut by linear cliquewidth). However, to the best of our knowledge, this is the first graph drawing (or computational geometry) problem shown to be XNLP-complete. The algorithms are described in Section 2, whereas the hardness is shown in Section 3. In contrast to the fact that OLP parameterized by the height lies in XP, it is not difficult to see that the socket/plug gadget described by Brückner and Rutter [7] can be utilized in the context of a reduction from 3-Partition to show that (PLP and, thus) CLP remains NP-hard even when restricted to instances of constant height. In fact, the unpublished full version of [7] features such a construction with a height of 7 [26]. Here, we present a reduction that is tailor-made for CLP, showing that it is NP-hard even when restricted to instances of height 4. We complement this result by showing that the (surprisingly challenging) case of instances with height at most three can be solved in polynomial time.
Theorem 1.2.
Constrained Level Planarity is NP-hard even when restricted to height 4, but instances of height at most 3 can be solved constructively in polynomial time.
We show the hardness in Appendix C, describe the algorithm in Section 4, and state open problems in Section 5. Proofs of statements marked with a (clickable) $\star$ are in the appendix.
@@ 2 An XP / XNLP Algorithm for Ordered Level Planarity
In this section, we show that Ordered Level Planarity is in XNLP (and thus in XP) when parameterized by the height $h$ of the input graph. Moreover, we show how to construct an ordered level planar drawing (if it exists) in XP-time. The main idea of our approach is to continuously sweep the plane with an unbounded y-monotone curve $s$ from left to right in a monotone fashion such that for each edge $(u,v)$, there is a point in time where $u$ and $v$ are consecutive vertices along $s$. When this happens, the edge can be drawn without introducing any crossings due to the fact that $s$ moves monotonically. To discretize this idea and turn it into an algorithm, we instead determine a sequence of unbounded y-monotone curves $S=(s_{1},s_{2},\dots,s_{z})$ that is sorted from left to right (i.e., no point of $s_{i+1}$ is to the left of $s_{i}$), has a length of $z\in\mathcal{O}(n)$, and contains for every edge $(u,v)$ a curve $s_{i}$ along which $u$ and $v$ are consecutive vertices. Now, given $S$, the desired drawing can be constructed in polynomial time; for an illustration see Fig. 1. Moreover, the sequence $S$ can be obtained in XP-time/space by exhaustively enumerating all possibilities or in XNLP-time/space by nondeterministic guessing. Let us proceed to formalize these ideas.
@@
A separation for $\mathcal{G}$ is an element of $P_{1}\times P_{2}\times\dots\times P_{h}$. Intuitively, a separation $s=(p_{1},p_{2},\dots,p_{h})$ represents the equivalence class of unbounded y-monotone curves that intersect line $L_{i}$ in position $p_{i}$ for each $1\leq i\leq h$; see Fig. 1. We say that a vertex $v\in V$ is on $s$ if $p_{\gamma(v)}$ represents $v$. Moreover, we say that $v$ is to the left (right) of $s$ if the index corresponding to the position of $v$ is strictly smaller (larger) than $p_{\gamma(v)}$. Consider two vertices $u,v\in V$ that are on $s$ and where $\gamma(u)<\gamma(v)$. We say that $u$ and $v$ are consecutive on $s$ if all of the indices $p_{\gamma(u)+1},p_{\gamma(u)+2},\dots,p_{\gamma(v)-1}$ represent gaps. In this case, $v$ is the successor of $u$ on $s$ and $u$ is the predecessor of $v$ on $s$. We say that $s$ uses an edge $e=(u,v)\in E$ if $u$ and $v$ are consecutive along $s$.
@@
[…]
@@
Let $S=(s_{1},s_{2},\dots,s_{z})$ be a sweeping sequence for our ordered level graph $\mathcal{G}$. We say that $S$ is nice if for each $i\in[z-1]$, the separation $s_{i+1}$ is obtained from $s_{i}$ by incrementing exactly one component by $1$. Moreover, we say that $S$ is exhaustive if it is nice and $s_{1}=(0,0,\dots,0)$ and $s_{z}=(|P_{1}|-1,|P_{2}|-1,\dots,|P_{h}|-1)$; see Fig. 1.
[…]
@@
Let $\mathcal{G}=(G=(V,E),\gamma,(\prec_{i})_{1\leq i\leq h})$ be an ordered level graph. Then $\mathcal{G}$ admits an ordered level planar drawing if and only if there is an exhaustive sweeping sequence for $G$ that uses every edge in $E$.
Proof 2.3 (Proof (sketch)).
The “if”-direction follows from Lemma 2.1. For the “only if”-direction, we show that the drawing can be internally triangulated by inserting y-monotone edges (allowing the use of parallel edges). For each y-monotone path $P$ that spans all levels, there is a nice sweeping sequence from $s_{1}=(0,\dots,0)$ to the separation corresponding to $P$ that uses all edges left of $P$. This is shown by arguing inductively on the number of triangles left of $P$.
@@
We remark that an exhaustive sweeping sequence using all edges corresponds directly to a particularly well-structured path decomposition. Thus, Lemma 2.2 implies that every (ordered) level planar drawing of height $h$ represents a graph of pathwidth at most $h-1$; a statement that was independently proven in [11]. However, the path decompositions constructed in the proof of [11, Lemma 1] do not exhibit the same properties that are inherent to exhaustive sweeping sequences and on which our algorithms heavily rely. In particular, these path decompositions may contain bags with multiple vertices of a given level (unless the drawing is proper). Moreover, the existence of a path decomposition of width at most $h-1$ for an ordered level graph $\mathcal{G}$ of height $h$ does not characterize the fact that $\mathcal{G}$ is ordered level planar (recall that, in fact, OLP is NP-hard even when restricted to instances of pathwidth $1$ [24]).
Figure 1: An exhaustive sweeping sequence using all edges of the depicted graph and the corresponding (cf. Lemma 2.1) ordered level planar drawing, as well as the corresponding sequence of true dynamic programming table entries $T[s,U]$ (cf. Lemma 2.4). Note that in any ordered level planar drawing of the graph, exactly one of the edges $(2,4),(1,3)$ is located to the left of the path $(1,2,3,4)$, while the other is located to the right. Similarly, in any exhaustive sweeping sequence containing separation $s_{5}$, exactly one of the edges $(2,4),(1,3)$ is used by a separation preceding $s_{5}$, while the other is used by a separation succeeding $s_{5}$. Hence, when iteratively building an exhaustive sweeping sequence, it is key to remember which edges between vertices of the current separation have already been used – this is exactly the purpose of the sets $U$. E.g., the fact that $(2,4)\in U_{5}$ corresponds to $(2,4)$ being used before $s_{5}$ (in $s_{4}$), from which one can infer how to proceed.
[…]
@@
Proof 2.5 (Proof (sketch)).
Let $\mathcal{G}=(G=(V,E),\gamma,(\prec_{i})_{1\leq i\leq h})$ be an ordered level graph and let $s=(p_{1},p_{2},\dots,p_{h})$ be a separation for $\mathcal{G}$. Further, let $U\subseteq E$ be a subset of the edges that are joining pairs of (not necessarily consecutive) vertices on $s$. We define $T[s,U]=\texttt{true}$ if there exists a nice sweeping sequence for $\mathcal{G}$ that starts with $s_{1}=(0,0,\dots,0)$, ends with $s$, and uses all edges in $U$, as well as all edges in $E$ incident to at least one vertex to the left of $s$. Otherwise, $T[s,U]=\texttt{false}$. Additionally, we allow $U=\bot$, in which case $T[s,U]=\texttt{false}$. Fig. 1 illustrates several true table entries $T[s_{i},U_{i}]$ along with the corresponding sweeping sequences $s_{1},s_{2},\dots,s_{i}$. Our goal is to determine $T[(|P_{1}|-1,|P_{2}|-1,\dots,|P_{h}|-1),\emptyset]$, which is true if and only if there exists an exhaustive sweeping sequence for $\mathcal{G}$ using all edges in $E$.
We will determine the entries $T[s,U]$ by means of a dynamic programming recurrence. For the base case, we simply set $T[(0,0,\dots,0),\emptyset]=\texttt{true}$. Now assume that $s\neq(0,0,\dots,0)$. For each index $1\leq j\leq h$ where $p_{j}\geq 1$, we define a separation
@@
For illustrations, refer to Fig. 1 (where for each $i\in[9]$ and $s=s_{i}$ and $U=U_{i}$, we have $s_{j}^{\prime}=s_{i-1}$ and $U_{j}^{\prime}=U_{i-1}$ for some $j\in[4]$). The following technical claim describes our recurrence relation.
[…]

### v2

@@ 1 Introduction
Klemz and Rote [27] showed that OLP (and, thus, CLP) is NP-hard even when restricted to the case where the underlying undirected graph of $G$ is a disjoint union of paths. Note that such a graph has bounded pathwidth (and treewidth), maximum degree, and feedback vertex set number, ruling out efficient parameterized algorithms with respect to these classical parameters. Additionally, the instances produced by their reduction have a levelwidth of only two. Independently, Brückner and Rutter [8] provided a very different reduction for showing that CLP is NP-hard (in fact, their reduction shows the NP-hardness of Partial Level Planarity, which can be seen as a generalization of OLP and a special case of CLP; see below). The instances constructed by their reduction are connected and have bounded maximum degree. They also present a polynomial time algorithm for CLP for the case where the graph has a single sink, which they later [9] sped up for the special case where the single-sink graph is also biconnected. Very recently, Klemz and Sieper [28] showed that CLP (and, thus, OLP) is FPT if parameterized by the vertex cover number of the input graph. They also observed that the reduction by Klemz and Rote (Brückner and Rutter) can easily be modified to show that OLP (CLP) is NP-hard even if restricting to (proper) instances of bounded treedepth.
@@
As discussed above, the parameterized complexity of OLP and CLP with respect to classical graph parameters (vertex cover number, feedback vertex set number, treedepth, pathwidth, treewidth, maximum degree) has been explored exhaustively [27, 8, 28]. In terms of more problem specific parameters, Klemz and Rote [27] showed that OLP (and, thus, CLP) is NP-hard even if restricted to instances of levelwidth two. In this paper, we focus on the other “dimension”: we study CLP and OLP parameterized by height.
@@
Parameterizing OLP by height captures the “linear” nature of the solution – this is reminiscent of recent results by Bodlaender, Groenland, Jacob, Jaffke, and Lima [6] who established XNLP-completeness for several problems parameterized by linear width measures (e.g., Capacitated Dominating Set by pathwidth and Max Cut by linear cliquewidth). However, to the best of our knowledge, this is the first graph drawing (or computational geometry) problem shown to be XNLP-complete. The algorithms are described in Section 2 (Theorem 3), whereas the hardness is shown in Section 3 (Theorem 5).
In contrast to the fact that OLP parameterized by the height lies in XP, it is not difficult to see that the socket/plug gadget described by Brückner and Rutter [8] can be utilized in the context of a reduction from 3-Partition to show that (PLP and, thus) CLP remains NP-hard even when restricted to instances of constant height. In fact, the unpublished full version of [8] features such a construction with a height of 7 [31]. Here, we present a reduction that is tailor-made for CLP, showing that it is NP-hard even if restricted to instances of height 4. We complement this result by showing that the (surprisingly challenging) case of instances with height at most 3 can be solved in polynomial time.
Theorem 2.
Constrained Level Planarity is NP-hard even if restricted to height 4, but instances of height at most 3 can be solved constructively in polynomial time.
We show the hardness in Section 4 (Theorem 6) and present the algorithm in Section 6 (Theorem 8). As a warm-up for the rather technical positive result, we show that CLP can be solved in linear time when restricted to instances of height at most 2; see Section 5 (Theorem 7).
We conclude by stating some open problems in Section 7.
@@ 2 An XP / XNLP Algorithm for Ordered Level Planarity
In this section, we show that Ordered Level Planarity is in XNLP (and thus in XP) if parameterized by the height $h$ of the input graph. Moreover, we show how to construct an ordered level planar drawing (if it exists) in XP-time. The main idea of our approach is to continuously sweep the plane with an unbounded y-monotone curve $s$ from left to right in a monotone fashion such that for each edge $(u,v)$, there is a point in time where $u$ and $v$ are consecutive vertices along $s$. When this happens, the edge can be drawn without introducing any crossings due to the fact that $s$ moves monotonically. To discretize this idea and turn it into an algorithm, we instead determine a sequence of unbounded y-monotone curves $S=(s_{1},s_{2},\dots,s_{z})$ that is sorted from left to right (i.e., no point of $s_{i+1}$ is to the left of $s_{i}$), has a length of $z\in\mathcal{O}(n)$, and contains for every edge $(u,v)$ a curve $s_{i}$ along which $u$ and $v$ are consecutive vertices. Now, given $S$, the desired drawing can be constructed in polynomial time; for an illustration see Figure 1. Moreover, the sequence $S$ can be obtained in XP-time/space by exhaustively enumerating all possibilities or in XNLP-time/space by nondeterministic guessing. Let us proceed to formalize these ideas.
@@
A separation for $\mathcal{G}$ is an element of $P_{1}\times P_{2}\times\dots\times P_{h}$. Intuitively, a separation $s=(p_{1},p_{2},\dots,p_{h})$ represents the equivalence class of unbounded y-monotone curves that intersect line $L_{i}$ in position $p_{i}$ for each $1\leq i\leq h$; see Figure 1. We say that a vertex $v\in V$ is on $s$ if $p_{\gamma(v)}$ represents $v$. Moreover, we say that $v$ is to the left (right) of $s$ if the index corresponding to the position of $v$ is strictly smaller (larger) than $p_{\gamma(v)}$. Consider two vertices $u,v\in V$ that are on $s$ and where $\gamma(u)<\gamma(v)$. We say that $u$ and $v$ are consecutive on $s$ if all of the indices $p_{\gamma(u)+1},p_{\gamma(u)+2},\dots,p_{\gamma(v)-1}$ represent gaps. In this case, $v$ is the successor of $u$ on $s$ and $u$ is the predecessor of $v$ on $s$. We say that $s$ uses an edge $e=(u,v)\in E$ if $u$ and $v$ are consecutive along $s$.
@@
Drawing algorithm.
The following lemma formalizes the idea of our drawing algorithm.
Lemma 1.
@@
Proof.
Without loss of generality, we assume that $s_{1}=(0,0,\dots,0)$ (otherwise, we simply prepend this separation to $S$). Recall that a separation $s_{j}=(p_{1},p_{2},\dots,p_{h})$ corresponds to an equivalence class of unbounded y-monotone curves that intersect line $L_{i}$ in position $p_{i}$ for each $1\leq i\leq h$. For each $j\in[z]$, let $c_{j}$ denote a member of the class of curves corresponding to $s_{j}$. The fact that $S$ is a sweeping sequence implies that we may assume that for every $j\in[z-1]$, no point of $c_{j+1}$ lies to the left of $c_{j}$ and that the only points shared by $c_{j}$ and $c_{j+1}$ are those that correspond to vertices. In particular, the curve $c_{1}$ passes through the left-most gap on each level.
To prove the existence of the desired drawing, it suffices to show that for each $j\in[z]$ there is an ordered level planar drawing $\Gamma_{j}$ that contains all vertices that are not to the right of $s_{j}$ and all edges that are used by one of the separations $s_{1},s_{2},\dots,s_{j}$ and where the vertices on $s_{j}$ and the edges used by $s_{j}$ are drawn on the curve $c_{j}$ and the remaining vertices and edges are drawn to the left of $c_{j}$. Note that, indeed, $\Gamma_{z}$ is the desired drawing of $\mathcal{G}$.
The (empty) drawing $\Gamma_{1}$ is trivially constructed. To construct $\Gamma_{j+1}$ from $\Gamma_{j}$, we simply place all (isolated) vertices that are to the right of $s_{j}$ and to the left of $s_{j+1}$ between $c_{j}$ and $c_{j+1}$ and then draw on $c_{j+1}$ all the new vertices on $s_{j+1}$, as well as the not yet drawn edges that are used by $s_{j+1}$. By definition of the curves $c_{1},c_{2},\dots,c_{j+1}$, this process cannot introduce any crossings. For illustrations, refer to Figure 1.
The above inductive proof directly corresponds to an iterative algorithm (note that we do not have to explicitly construct the curves $c_{1},c_{2},\dots,c_{z}$). The total time for adding the vertices is $\mathcal{O}(zh+n)$. To add the edges, whenever there is a new pair of consecutive vertices on the current separation, we need to check whether these vertices are adjacent, which takes $\mathcal{O}(n)$ time. New pairs of consecutive vertices are only created when a vertex appears or disappears from the current separation. Moreover, the appearance/disappearance of a vertex can only create $\mathcal{O}(1)$ new pairs of consecutive vertices. Combined with the fact that for each vertex there is an interval of separations in $S$ on which it is located, it follows that the total number of adjacency checks we have to perform is $\mathcal{O}(n)$. Thus, the total time for the adjacency checks is $\mathcal{O}(n^{2})$. Drawing an edge takes time linear in the number of levels it spans, which is at most $h\in\mathcal{O}(n)$. Moreover, by planarity (note that the inductive proof shows that $G$ is planar), the number of edges is $\mathcal{O}(n)$. Hence, the time to draw all the edges is $\mathcal{O}(n^{2})$, and the total runtime is $\mathcal{O}(zh+n^{2})$, as claimed. ∎
In the remainder of this section, we will show that for each ordered level graph that admits an ordered level planar drawing there exists a particularly well-structured sweeping sequence that uses all of its edges. Moreover, we will show that the existence of such a sequence can be tested efficiently. Combined with Lemma 1, this results in the desired XP / XNLP algorithms.
Nice and exhaustive sweeping sequences.
@@
Let $S=(s_{1},s_{2},\dots,s_{z})$ be a sweeping sequence for our ordered level graph $\mathcal{G}$. We say that $S$ is nice if for each $i\in[z-1]$, the separation $s_{i+1}$ is obtained from $s_{i}$ by incrementing exactly one component by $1$. Moreover, we say that $S$ is exhaustive if it is nice and $s_{1}=(0,0,\dots,0)$ and $s_{z}=(|P_{1}|-1,|P_{2}|-1,\dots,|P_{h}|-1)$; see Figure 1.
Lemma 2.
@@
Let $\mathcal{G}=(G=(V,E),\gamma,(\prec_{i})_{1\leq i\leq h})$ be an ordered level graph. Then $\mathcal{G}$ admits an ordered level planar drawing if and only if there is an exhaustive sweeping sequence for $\mathcal{G}$ that uses every edge in $E$.
Proof.
The “if”-direction follows immediately from Lemma 1. For the “only if”-direction, suppose that $\mathcal{G}$ admits an ordered level planar drawing $\Gamma$. We augment $\mathcal{G}$ and $\Gamma$ as follows: first, we insert a new bottommost and topmost level with a single vertex $a$ and $b$, respectively. Second, we insert a maximal set of y-monotone crossing-free edges (i.e., the resulting drawing is supposed to be crossing-free). We explicitly allow the introduction of parallel edges, but we require that for each pair of parallel edges $e,e^{\prime}$ there is a vertex in the interior of the simple closed curve formed by $e,e^{\prime}$. Let $\mathcal{G}^{\prime}$ and $\Gamma^{\prime}$ denote the resulting graph and drawing, respectively, which is an internal triangulation:
Claim 1.
The outer face of $\Gamma^{\prime}$ is bounded by two parallel $(a,b)$-edges and every internal face of $\Gamma^{\prime}$ is a triangle (i.e., its boundary corresponds to a cycle of length three).
Proof.
The statement regarding the outer face obviously holds. Now consider an internal face $f$. Towards a contradiction, assume that $f$ is not a triangle. Let $v$ be a bottommost vertex of $f$, which is incident to (at least) two edges of $f$. Let $e_{1}=(v,v_{1}),e_{2}=(v,v_{2})$ be two edges that are consecutive in the cyclic order of edges incident to $v$ and consecutive along the boundary of $f$. Without loss of generality, we may assume that $\gamma(v_{2})\leq\gamma(v_{1})$. Moreover, without loss of generality, we may assume that $e_{2}$ is to the right of $e_{1}$. We distinguish two main cases.
Case 1.  There is a vertex $u$ with $\gamma(v)<\gamma(u)\leq\gamma(v_{2})$ that is located to the right of $e_{1}$ and to the left of $e_{2}$. We can add an edge from $v$ to the bottommost vertex with these properties in a y-monotone crossing-free fashion in $f$. Moreover, this edge cannot be parallel to some edge of $\Gamma^{\prime}$; a contradiction to the definition of $\Gamma^{\prime}$.
Case 2.  There is no vertex $u$ with $\gamma(v)<\gamma(u)\leq\gamma(v_{2})$ that is located to the right of $e_{1}$ and to the left of $e_{2}$. We distinguish two subcases.
Case 2.1.$\gamma(v_{2})<\gamma(v_{1})$. In this case, we can add an edge from $v_{2}$ to $v_{1}$ in a y-monotone crossing-free fashion in $f$ such that, even if this edge is parallel to some other edge, the interior of the region bounded by these edges contains at least one vertex (recall that $f$ is not a triangle by assumption). So again, we obtain a contradiction to the definition of $\Gamma^{\prime}$.
Case 2.2.$\gamma(v_{2})=\gamma(v_{1})$. In this case, we draw a y-monotone curve $c$ in $f$ that starts at $v$ and extends upwards until it crosses some edge $e_{3}=(u,w)$ with $e_{3}\notin\{e_{1},e_{2}\}$. The existence of $e_{3}$ follows from the existence of the two edges of the outer face. By applying the assumption of Case 2 to $u$, it follows that the upper endpoint $w$ of $e_{3}$ is neither $v_{1}$ nor $v_{2}$. Hence, we can add an edge from $v$ to $w$ in a y-monotone crossing-free fashion in $f$ such that, even if this edge is parallel to some other edge, the interior of the region bounded by these edges contains at least one vertex ($v_{1}$ or $v_{2}$). So again, we obtain a contradiction to the definition of $\Gamma^{\prime}$. ∎
Let $e_{\ell}$ and $e_{r}$ denote the left and right edge of the outer face of $\Gamma^{\prime}$. To prove the lemma, it suffices to show that there is a nice sweeping sequence for $\mathcal{G}^{\prime}$ that uses all of its edges and starts with separation $s_{1}=(1,0,0,\dots,1)$ and ends with separation $s_{z}=(1,|P_{1}|-1,|P_{2}|-1,\dots,|P_{z}|-1,1)$, which use $e_{\ell}$ and $e_{r}$, respectively. To this end, we show:
Claim 2.
Let $P$ be a directed (y-monotone) $ab$-path in $\Gamma^{\prime}$ and let $s$ be the (unique) separation that corresponds to $P$. Then there is a nice sweeping sequence from $s_{1}$ to $s$ that uses all edges of $\Gamma^{\prime}$ that are not located to the right of $P$.
Proof.
The proof is by induction on the number $i$ of triangles to the left of $P$. If $i=0$ (that is, $P=(e_{\ell})$), the statement obviously holds. So assume $i>0$.
Suppose there is an edge $(u,v)$ on $P$ that is incident to a triangle $uvw$, where $w$ is to the left of $P$ and $\gamma(u)<\gamma(w)<\gamma(v)$; we call this Configuration (a). In this case, we can replace $(u,v)$ with $(u,w)$ and $(w,v)$ to obtain a directed $ab$-path $P^{\prime}$ from $P$ with $i-1$ triangles to its left and the claim follows by induction.
Next, suppose that there is a triangle $uvw$ to the left of $P$ whose edges $(u,w)$ and $(w,v)$ belong to $P$; we call this Configuration (b). We can replace $(u,w)$ and $(w,v)$ with $(u,v)$ to obtain a directed $ab$-path $P^{\prime}$ from $P$ with $i-1$ triangles to its left and the claim follows by induction.
We claim that one of these two configurations always exists. Towards a contradiction, assume otherwise. Let $e=(u,v)$ be an edge on $P$ that is incident to a triangle $\triangle=uvw$, where $w$ is to the left of $P$. Since Configuration (a) does not exist, we may assume without loss of generality that $(\gamma(u)<)\gamma(v)<\gamma(w)$. Further, we may assume without loss of generality that $e$ is the topmost edge on $P$ with this property. Since Configuration (b) does not exist, the edge $(v,w)$ of $\triangle$ does not belong to $P$. Let $x$ denote the successor of $v$ along $P$ and let $vxy$ denote the triangle to the left of $(v,x)$ on $P$. Since the edge $(v,w)$ of $\triangle$ is crossing-free and Configuration (b) does not exist, it follows that $(\gamma(v)<)\gamma(x)<\gamma(y)$; a contradiction to the choice of $e$. ∎
The desired sequence from $s_{1}$ to $s_{z}$ is obtained by Claim 2, choosing $P=(e_{r})$. ∎
@@
We remark that an exhaustive sweeping sequence using all edges corresponds directly to a particularly well-structured path decomposition. Thus, Lemma 2 implies that every (ordered) level planar drawing of height $h$ represents a graph of pathwidth at most $h-1$; a statement that was independently proven in [12]. However, the path decompositions constructed in the proof of [12, Lemma 1] do not exhibit the same properties that are inherent to exhaustive sweeping sequences and on which our algorithms heavily rely. In particular, these path decompositions may contain bags with multiple vertices of a given level (unless the drawing is proper). Moreover, the existence of a path decomposition of width at most $h-1$ for an ordered level graph $\mathcal{G}$ of height $h$ does not characterize the fact that $\mathcal{G}$ is ordered level planar (recall that, in fact, OLP is NP-hard even when restricted to instances of pathwidth $1$ [27]).
Computing suitable sweeping sequences.
In view of Lemma 2, recognizing ordered level planar graphs is equivalent to testing for the existence of suitable sweeping sequences. We now show that this can be done efficiently.
Lemma 3.
@@
Proof.
Let $\mathcal{G}=(G=(V,E),\gamma,(\prec_{i})_{1\leq i\leq h})$ be an ordered level graph, let $s=(p_{1},p_{2},\dots,p_{h})$ be a separation for $\mathcal{G}$, and let $U\subseteq E$ be a subset of the edges that are joining pairs of (not necessarily consecutive) vertices on $s$. We define $T[s,U]=\texttt{true}$ if there exists a nice sweeping sequence for $\mathcal{G}$ that starts with $s_{1}=(0,0,\dots,0)$, ends with $s$, and uses all edges in $U$, as well as all edges in $E$ incident to at least one vertex to the left of $s$. Otherwise, $T[s,U]=\texttt{false}$. Additionally, we allow $U=\bot$, in which case $T[s,U]=\texttt{false}$. Figure 1 illustrates several true table entries $T[s_{i},U_{i}]$ along with the corresponding sweeping sequences $s_{1},s_{2},\dots,s_{i}$. Our goal is to determine $T[(|P_{1}|-1,|P_{2}|-1,\dots,|P_{h}|-1),\emptyset]$, which is true if and only if there exists an exhaustive sweeping sequence for $\mathcal{G}$ using all edges in $E$.
Recurrence relation.  We will determine the entries $T[s,U]$ by means of a dynamic programming recurrence. For the base case, we simply set $T[(0,0,\dots,0),\emptyset]=\texttt{true}$. Now assume that $s\neq(0,0,\dots,0)$. For each index $1\leq j\leq h$ where $p_{j}\geq 1$, we define a separation
@@
Figure 1: An exhaustive sweeping sequence using all edges of the depicted graph and the corresponding (cf. Lemma 1) ordered level planar drawing, as well as the corresponding sequence of true dynamic programming table entries $T[s,U]$ (cf. Lemma 3). Note that in any ordered level planar drawing of the graph, exactly one of the edges $(2,4),(1,3)$ is located to the left of the path $(1,2,3,4)$, while the other is located to the right. Similarly, in any exhaustive sweeping sequence containing separation $s_{5}$, exactly one of the edges $(2,4),(1,3)$ is used by a separation preceding $s_{5}$, while the other is used by a separation succeeding $s_{5}$. Hence, when iteratively building an exhaustive sweeping sequence, it is key to remember which edges between vertices of the current separation have already been used – this is exactly the purpose of the sets $U$. E.g., the fact that $(2,4)\in U_{5}$ corresponds to $(2,4)$ being used before $s_{5}$ (in $s_{4}$), from which one can infer how to proceed.
For illustrations, refer to Figure 1 (where for each $i\in[9]$ and $s=s_{i}$ and $U=U_{i}$, we have $s_{j}^{\prime}=s_{i-1}$ and $U_{j}^{\prime}=U_{i-1}$ for some $j\in[4]$). The following technical claim describes our recurrence relation.
Claim 3.